1.3) (15 points) Exercise 2.1.18. Except there are 106 parts and the numbers in the first column change from 70, 12 to 81, 11 respectively. Discs of polycarbonate plastic from a supplier are analyzed for scratch and shock resistance. The results from 106 discs are summarized as follows: Shock Resistance High 81 Low Scratch High Resistance Low Let A denote the event that a sample has high shock resistance, and let B denote the event that a disc has high scratch resistance. If a disc is selected at random, determine the number of samples in A nB, A and A UB. stho nasswerds contain
Answer #1

Let A denote the event that a sample has high shock resistance, and B denote the event that a disc has high scratch resistance.

A' denote the event that a sample has low shock resistance and B' denote the event that a disc has low scratch resistance.

Therefore, the number of samples in following events are

=> n ( A cap B) = 81

n(A)9+5 14

n(A) = 81 + 11 = 92

n(B) = 81 + 9 = 90

Therefore,

n(A U B) n(A) + n(B) _ n(A n B)-92+ 90-81 => n(AU B) = 101